# Quantum theory of surface lattice resonances

### v1

@@ II Application 1: Molecular optomechanics with SLRs
III Application 2: Nonlinear switching of excitonic SLRs
So far, the scope of this paper has been devoted to treating the nanoparticles providing the SLRs in the linear optical regime. An interesting generalization of the framework is given by the extension to the nonlinear optical regime, which can, e.g., stem from anharmonicities of the particles. One possible realization of this consists in replacing the MNPs with a molecular material featuring tunable excitonic resonances such as proposed in Ref. Humphrey et al. (2016).
Figure 5: (a) Nonlinear SLR with array of saturable emitters: Pumping excitation into the |$\ket{2}$ state shuts off the polarizability of the |$\ket{1}\to\ket{2}$ transition while opening up the dipole transition from |$\ket{2}\to\ket{3}$ whose frequency matches the first diffractive mode of the lattice. (b) Extinction spectrum of an array of $M=10^{3}$ three-level systems if all of the population is in the ground state (top) and if half of the population is pumped to the |$\ket{2}$ state (bottom). The lattice spacing is $a=415\>\mathrm{nm}$. We chose the transition frequencies $\hbar\omega_{12}\approx 1.5\>\mathrm{eV},\,\hbar\omega_{13}\approx 3.0\>\mathrm{eV}$, as well as the radiative linewidths $\Gamma_{12}^{\mathrm{rad}}=\Gamma_{23}^{\mathrm{rad}}=0.25\>\mathrm{eV}$. The $\omega_{12}$ and $\omega_{23}$ transitions corresponds to a resonance wavelengths of $\lambda_{12}=800\>\mathrm{nm}$ and $\lambda_{23}=400\>\mathrm{nm}$, respectively. The dotted white lines show the frequencies corresponding to the Rayleigh anomalies. The vertical dashed green lines mark cross sections through the extinction spectrum at $k_{\parallel}=0$ which are shown in (c).
In the following, we will show that the nonlinearity of electronic transitions can be used to switch the SLR condition. This is enabled by ‘switching off’ a transition that does not fulfill the Rayleigh anomaly condition and ‘switching on’ a transition matching the Rayleigh criterion for transition between two excited states. In particular, we consider a chain of anharmonic three-level chromophores described by states |$\ket{\nu}_{j}$ where $j$ is the index describing the position of the particle in the lattice and $\nu$ describes the internal electronic label at energies $\hbar\omega_{\nu}$, $\nu=1,2,3$. These could for instance correspond the $S_{0}$, $S_{1}$ and $S_{2}$ states of an organic molecule. The transition dipole operator for a single molecule $j$ can then be expressed as $\hat{\boldsymbol{\mu}}^{j}=\sum_{\nu<\nu^{\prime}}\hat{\boldsymbol{\mu}}_{\nu\nu^{\prime}}^{j}$ where $\hat{\boldsymbol{\mu}}_{\nu\nu^{\prime}}^{j}=\boldsymbol{\mu}_{\nu\nu^{\prime}}(\hat{\sigma}_{j}^{\nu\nu^{\prime}}+\mathrm{H.c.})$, with the operator for the $\nu\nu^{\prime}$ transition $\hat{\sigma}_{j}^{\nu\nu^{\prime}}=\ket{\nu}_{j}\bra{\nu^{\prime}}_{j}$, and the dipole moment between two transitions is computed as $\boldsymbol{\mu}_{\nu\nu^{\prime}}=\bra{\nu}_{j}\hat{\boldsymbol{\mu}}_{\nu\nu^{\prime}}^{j}\ket{\nu^{\prime}}_{j}$ which we assume to be identical for all molecules. We furthermore assume that all transition dipoles are oriented along the same direction. The free Hamiltonian of the excitonic lattice can then be expressed as
^$\hat{\mathcal{H}}_{\text{ex}}=\hbar\sum_{\nu=1}^{3}\omega_{\nu}\sum_{j=1}^{M}\ket{\nu}_{j}\bra{\nu}_{j}.$ (16)
One can then derive an expression for the effective polarizability of the $\nu\nu^{\prime}$ transition (for details see SM) Mukamel (1995)
$\boldsymbol{\alpha}_{\mathbf{q}}^{\mathrm{eff}}(\omega)_{\nu\nu^{\prime}}=-\frac{\mathrm{i}p_{\text{inv}}^{\nu\nu^{\prime}}(\boldsymbol{\mu}_{\nu\nu^{\prime}}\cdot\boldsymbol{\mu}_{\nu\nu^{\prime}}^{\top})}{\left[\Gamma_{\nu\nu^{\prime}}^{\mathrm{rad}}/2-\mathrm{i}(\omega-\omega_{\nu\nu^{\prime}})+\mathrm{i}\mathcal{S}_{\mathbf{q}}(\omega)^{\nu\nu^{\prime}}p_{\text{inv}}^{\nu\nu^{\prime}}\right]},$ (17)
@@ III Application 2: Nonlinear switching of excitonic SLRs
For illustrative purposes, let us consider a simplified scenario in the following where only the $1\to 2$ and $2\to 3$ transitions possess a non-vanishing dipole moment, i.e., $\mu_{13}=0$. The corresponding extinction spectra are plotted in Fig. 5(b) while cross sections through the extinction spectrum at $|\mathbf{k}_{\parallel}|=0$ are shown in Fig. 5(c). Here, we choose a scenario where the $\omega_{12}$ transition does not match the frequency corresponding to the Rayleigh criterion (around $|\mathbf{k}_{\parallel}|=0$). The extinction spectrum on top of Fig. 5(c) shows a modification of the linewidth as well as the resonance frequency as opposed to a bare molecule due to the dipolar interactions in the lattice but there is no SLR. On the other hand, we chose the $\omega_{23}$ transition such that it matches the frequency of the $m=1$ Rayleigh anomaly modes at $|\mathbf{k}_{\parallel}|=0$. The resulting extinction spectrum at $|\mathbf{k}_{\parallel}|=0$ now shows much stronger modification as opposed to the bare molecule spectrum, with a narrow SLR feature as discussed in Sec. I.
IV Conclusions, outlook and extensions
@@
In the first application, we discussed how high-$Q$ SLRs can providing a pathway to bring molecular optomechanics in the resolved sideband regime and mediated the coupling between nanoparticle arrays and collective vibrational modes. In the second application, we demonstrated the versatility of SLRs in the nonlinear optical regime by considering arrays of multi-level emitters. Through optical pumping, the excitonic nonlinearity allowed for selective activation of specific SLR conditions corresponding to different electronic transitions. This ability to dynamically control SLRs suggests potential uses in tunable photonic devices and optically driven switches, particularly in systems where reconfigurable optical properties are desirable.
For future endeavors, the quantized description of the electromagnetic field employed in our treatment could e.g., be used to investigate non-classical states of light such as sequeezed light input, with possible applications in quantum-enhanced sensing. While so far, as a proof of principle, we have focused on extinction spectra, in principle we could also compute more complex material observables, like molecular populations through nonlinear optical spectroscopy, using mean-field approaches Li et al. (2018); Fowler-Wright et al. (2022); Reitz et al. (2024). Finally, while we have focused on a simplified treatment of the nanoparticles as harmonic oscillators in the dipole approximation, the approach can be extended to more complicated treatments, e.g., based on quasi-normal mode theory or higher-order multipoles Miwa and Schatz (2021).
V Acknowledgments
@@
×$\times\mathrm{e}^{-\mathrm{i}\mathbf{q}\cdot\mathbf{r}_{j}}\boldsymbol{\varepsilon}_{\boldsymbol{\mu}}\cdot\mathrm{Im}[\mathbf{G}(\mathbf{r}_{j},\mathbf{r}_{j^{\prime}},\omega_{\mathbf{k}})]\cdot\boldsymbol{\varepsilon}_{\boldsymbol{\mu}}\mathrm{e}^{\mathrm{i}\mathbf{q}^{\prime}\cdot\mathbf{r}_{j^{\prime}}}{\hat{A}}_{\mathbf{q}^{\prime}}(t^{\prime}).$
@@
×$\times\sum_{\Lambda}\mathrm{e}^{-\mathrm{i}\mathbf{q}\cdot\mathbf{r}_{\Lambda}}\boldsymbol{\varepsilon}_{\boldsymbol{\mu}}\cdot\mathrm{Im}\left[\mathbf{G}(\mathbf{r}_{\Lambda},\omega_{\mathbf{k}})\right]\cdot\boldsymbol{\varepsilon}_{\boldsymbol{\mu}}{\hat{A}}_{\mathbf{q}}(\omega)\,\mathrm{e}^{\mathrm{i}\omega(t-t^{\prime})}.$ (S18)
@@
in terms of the single-particle radiative rate $\Gamma_{0}^{\mathrm{rad}}=(\mu_{0}^{2}\omega_{0}^{3})/(3\pi c^{3}\hbar\epsilon_{0})$. Here, the reduced lattice sum $\mathcal{S}_{\mathbf{q}}(\omega)$ now excludes the yero-discplacement contribution and we made a Markovian assumption for the single-particle radiative rate only by evaluating it around $\omega_{0}\approx\omega$. The condition that needs to be fulfilled for SLRs is that both the real and the imaginary part of the expression in the brackets becomes (close to) zero. For numerical calculations, it is useful to reexpress the constants appearing in front of the lattice sum in terms of the single-particle radiative decay rate as
@@
with $\lambda(\omega)=2\pi c/\omega$.
@@
Figure S1: Illustration of a nanoparticle array ($\hat{\boldsymbol{\mu}}_{j}$) and an array of dipolar emitters ($\hat{\mathbf{d}}_{j}$) which are commonly coupled to the background electromagnetic field modes $\hat{a}_{\mathbf{k},\lambda}$. Elimination of the electromagnetic field leads to effective interactions both within the arrays and between the nanoparticle and emitter array, which are encompassed in the corresponding lattice sums.
Let us now consider the additional coupling of the nanoparticle array to an (e.g., molecular) array of dipolar emitters. We use this as an instructive case to facilitate the derivation of the optomechanical interaction in the following section. We assume the individual emitters with resonance frequency $\omega_{d}$ located at positions $\mathbf{r}_{j}^{\mathrm{m}}=\mathbf{r}_{j}+\mathbf{r}_{\mathrm{m}}$ described by the dipole operator $\hat{\mathbf{d}}_{j}=\mathbf{d}_{0}\left(\hat{B}_{j}+\hat{B}_{j}^{\dagger}\right)$ with dipole moment $\mathbf{d}_{0}$ and annihilation/creation operators $[\hat{B}_{j}^{\phantom{\dagger}},\hat{B}_{j^{\prime}}^{\dagger}]=\delta_{j,j^{\prime}}$ and typically $|\mathbf{d}_{0}|\ll|\boldsymbol{\mu}_{0}|$. We also assume that the emitter array has the same periodicity as the nanoparticle array. We start again by writing equations of motion for dipole operators as well as the electromagnetic field operators
@@
and replacing them in the equations for the dipole operators. The procedure is now largely identical to Sec. S1. Derivation of non-Markovian lattice response and consists in going to Fourier space both in the time and spatial domains. Finally, one ends up with an algebraic set of equations for the dipole operators
@@
describing a sum over all interactions between nanoparticle and emitter array (see Fig. S1 for pictorial representation of lattice sums). The input noise terms $\hat{A}_{\mathbf{q}}^{\mathrm{in}}(\omega),\hat{B}_{\mathbf{q}}^{\mathrm{in}}(\omega)$ are defined analogously to Sec. S1. Derivation of non-Markovian lattice response. Assuming only the nanoparticles to be driven by a non-zero average input electric field, i.e., $\expectationvalue{\hat{B}_{\mathrm{in}}(\omega)}=0$, the dipole amplitude of the nanoparticle array can be expressed as
@@
^$\hat{\mathcal{H}}_{\mathrm{OM}}=-\sum_{j}\hat{\mathbf{p}}_{j}^{R}\cdot\hat{\mathbf{E}}(\mathbf{r}_{j}^{\text{m}}).$ (S32)
@@
and the free energy of the vibrational modes is given by ^$\hat{\mathcal{H}}_{\mathrm{vib}}=\sum_{j}\omega_{\mathrm{vib}}\hat{b}_{j}^{\dagger}\hat{b}_{j}^{\phantom{\dagger}}$. With this, the optomechanical Hamiltonian can be written as
^$\hat{\mathcal{H}}_{\mathrm{OM}}=-\sum_{j}\sum_{\mathbf{k},\lambda}\mathcal{E}_{\mathbf{k}}\left(\mathbf{p}_{j}\mathrm{e}^{-\mathrm{i}\omega_{\ell}t}+\mathbf{p}_{j}^{*}\mathrm{e}^{\mathrm{i}\omega_{\ell}t}\right)(\hat{b}_{j}^{\dagger}+\hat{b}_{j}^{\phantom{\dagger}})\left(\boldsymbol{\varepsilon}_{\mathbf{k},\lambda}\hat{a}_{\mathbf{k},\lambda}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{m}}+\boldsymbol{\varepsilon}_{\mathbf{k},\lambda}^{*}\hat{a}_{\mathbf{k},\lambda}^{\dagger}\mathrm{e}^{-\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{m}}\right).$ (S35)
@@
^$\hat{\mathcal{H}}_{\mathrm{OM}}=\hbar\sum_{j}\sum_{\mathbf{k},\lambda}\left[g_{\mathbf{k},\lambda}^{\mathrm{OM}}\hat{a}_{\mathbf{k},\lambda}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{\mathrm{m}}}\mathrm{e}^{\mathrm{i}\omega_{\ell}t}+g_{\mathbf{k},\lambda}^{\mathrm{OM},*}\hat{a}_{\mathbf{k},\lambda}^{\dagger}\mathrm{e}^{-\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{\mathrm{m}}}\mathrm{e}^{-\mathrm{i}\omega_{\ell}t}\right](\hat{b}_{j}^{\dagger}+\hat{b}_{j}^{\phantom{\dagger}}),$ (S36)
@@
^$\hat{\mathcal{H}}_{\mathrm{OM}}^{\mathrm{AS}}\approx\hbar\sum_{j}\sum_{\mathbf{k},\lambda}\left[g_{\mathbf{k},\lambda}^{\mathrm{OM}}\hat{a}_{\mathbf{k},\lambda}\hat{b}_{j}^{\dagger}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{\mathrm{m}}}\mathrm{e}^{\mathrm{i}\omega_{\ell}t}+g_{\mathbf{k},\lambda}^{\mathrm{OM},*}\hat{a}_{\mathbf{k},\lambda}^{\dagger}\hat{b}_{j}\mathrm{e}^{-\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{\mathrm{m}}}\mathrm{e}^{-\mathrm{i}\omega_{\ell}t}\right],$ (S37)
@@
^$\hat{\mathcal{H}}_{\mathrm{OM}}^{\mathrm{S}}\approx\hbar\sum_{j}\sum_{\mathbf{k},\lambda}\left[g_{\mathbf{k},\lambda}^{\mathrm{OM}}\hat{a}_{\mathbf{k},\lambda}\hat{b}_{j}^{\phantom{\dagger}}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{\mathrm{m}}}\mathrm{e}^{\mathrm{i}\omega_{\ell}t}+g_{\mathbf{k},\lambda}^{\mathrm{OM},*}\hat{a}_{\mathbf{k},\lambda}^{\dagger}\hat{b}_{j}^{\dagger}\mathrm{e}^{-\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{\mathrm{m}}}\mathrm{e}^{-\mathrm{i}\omega_{\ell}t}\right],$ (S43)
@@
S4. Nonlinear polarizability of excitonic SLRs

### v2

@@ III Application 1: Molecular optomechanics with SLRs
IV Application 2: Nonlinear switching of excitonic SLRs
So far, the scope of this paper has been devoted to treating the nanoparticles providing the SLRs in the linear optical regime, corresponding to harmonic oscillators. An interesting generalization of the framework is given by the extension to the nonlinear optical regime, which can, e.g., stem from anharmonicities of the particles. One possible realization of this consists in replacing the MNPs with a molecular material featuring tunable excitonic resonances such as proposed in Ref. Humphrey et al. (2016).
@@ IV Application 2: Nonlinear switching of excitonic SLRs
[…]
One can then derive an expression for the effective polarizability of the $\nu\nu^{\prime}$ transition (which despite describing a signal that includes nonlinearities, can be computed in linear response) (for details see SM) Mukamel (1995)
$\boldsymbol{\alpha}_{\mathbf{q}}^{\mathrm{eff}}(\omega)_{\nu\nu^{\prime}}=-\frac{\mathrm{i}p_{\text{inv}}^{\nu\nu^{\prime}}(\boldsymbol{\mu}_{\nu\nu^{\prime}}\cdot\boldsymbol{\mu}_{\nu\nu^{\prime}}^{\top})}{\left[\mathrm{i}\mathcal{S}_{\mathbf{q}}^{\nu\nu^{\prime}}(\omega)p_{\text{inv}}^{\nu\nu^{\prime}}-\mathrm{i}(\omega-\omega_{\nu\nu^{\prime}})+\Gamma_{\nu\nu^{\prime}}^{\mathrm{rad}}/2\right]},$ (17)
[…]
@@
For illustrative purposes, let us consider a simplified scenario in the following where only the $1\to 2$ and $2\to 3$ transitions possess a non-vanishing dipole moment, i.e., $\mu_{13}=0$. The corresponding extinction spectra are plotted in Fig. 5(b) while cross sections through the extinction spectrum at $|\mathbf{k}_{\parallel}|=0$ are shown in Fig. 5(c). Here, we choose a scenario where the $\omega_{12}$ transition does not match the frequency corresponding to the Rayleigh criterion (around $|\mathbf{k}_{\parallel}|=0$). The extinction spectrum on top of Fig. 5(c) shows a modification of the linewidth as well as the resonance frequency as opposed to a bare molecule due to the dipolar interactions in the lattice but there is no SLR. On the other hand, we chose the $\omega_{23}$ transition such that it matches the frequency of the $m=1$ Rayleigh anomaly modes at $|\mathbf{k}_{\parallel}|=0$. The resulting extinction spectrum at $|\mathbf{k}_{\parallel}|=0$ now shows much stronger modification as opposed to the bare molecule spectrum, with a narrow SLR feature as discussed in the model section.
V Pump-probe spectroscopy of nonlinear excitonic SLR dynamics
In this section, we show how dynamical evolution of the populations can be taken into account by performing a perturbative expansion in the input field Mukamel (1995); Reitz et al. (2025). To this end, we model a pump–probe experiment with two impinging pulses, where a weak pump ($p$) prepares a nonequilibrium population distribution that is subsequently interrogated by a weak probe ($p^{\prime}$) [see Fig. 6(a)]. We assume the central frequency of the pump pulse to be (approximately) resonant with the $1\!\leftrightarrow\!2$ transition $\omega_{p}\approx\omega_{12}$, and the frequency of the probe pulse to be (approximately) resonant with the $2\!\leftrightarrow\!3$ transition $\omega_{p^{\prime}}\approx\omega_{23}$. Further, we assume the two transitions to be very off-resonant such that cross-couplings between them may be neglected. The (classical) input fields affecting the excitonic transitions of the $j$-th emitter can then be written as
$\langle\hat{\sigma}_{j}^{12,\mathrm{in}}(t)\rangle$ $=\eta_{p}f_{p}(t)\,\mathrm{e}^{-\mathrm{i}\omega_{p}t}\mathrm{e}^{\mathrm{i}\mathbf{k}_{\parallel}\cdot\mathbf{r}_{j}},$ (18a)
$\langle\hat{\sigma}^{23,\mathrm{in}}_{j}(t)\rangle$ $=\eta_{p^{\prime}}f_{p^{\prime}}(t-\tau_{\Delta})\,\mathrm{e}^{-\mathrm{i}\omega_{p^{\prime}}t}\mathrm{e}^{\mathrm{i}\mathbf{k}_{\parallel}\cdot\mathbf{r}_{j}},$ (18b)
where $\eta_{p,p^{\prime}}$ describes the pulse amplitude, $f_{p,p^{\prime}}(t)$ the temporal envelope (assumed Gaussian here), $\tau_{\Delta}$ is the delay time between the pulses, and we assume collinear incidence of pump and probe pulses at wavevector parallel to the array plane $\mathbf{k}_{\parallel}$. Following Refs. Mukamel (1995); Reitz et al. (2025), we now expand all system operators $\hat{O}_{\mathbf{q}}$ in terms of the pump and probe amplitudes (see Sec. S5 of the SM for details)
$\langle\hat{O}_{\mathbf{q}}\rangle=\sum_{n,m}\eta_{p}^{n}\eta_{p^{\prime}}^{m}\langle\hat{O}_{\mathbf{q}}\rangle^{(n)(m)}.$ (19)
Assuming factorizability between coherences and populations, this yields a linearized set of equations up to $(n),(m)$-th order. In the following, we are interested in the linear response to the pump and the linear response to the probe, where the probe interrogates pump-induced second-order populations, yielding an overall third-order correction. At first order, the pump can create coherence for the $1\!\leftrightarrow\!2$ transition:
$\langle\hat{\sigma}_{\mathbf{q}}^{12}(\omega)\rangle^{(1)(0)}=\frac{p_{\mathrm{inv}}^{12}(0)f_{p}(\omega-\omega_{p})\delta_{\mathbf{q},\mathbf{k}_{\parallel}}}{\mathrm{i}\mathcal{S}_{\mathbf{q}}^{12}(\omega)p_{\text{inv}}^{12}(0)-\mathrm{i}(\omega-\omega_{12})+\Gamma_{12}^{\mathrm{rad}}/2},$ (20)
[…]
$\langle\hat{\sigma}_{\mathbf{q}}^{23}(\omega)\rangle^{(2)(1)}_{\Delta\tau}=\frac{[f_{p^{\prime}}(\omega-\omega_{p^{\prime}})_{\Delta\tau}\ast\langle\hat{\sigma}^{12,\dagger}\hat{\sigma}^{12}\rangle_{\mathbf{q}-\mathbf{k}_{\parallel}}^{(2)(0)}](\omega)}{\Gamma_{23}^{\mathrm{rad}}/2-\mathrm{i}\mathcal{S}_{\mathbf{q}}^{23}(\omega)-\mathrm{i}(\omega-\omega_{23})},$ (21)
involving a convolution in frequency domain between the probe pulse and the second-order population created by the pump at zero quasi-momentum $\langle\hat{\sigma}^{12,\dagger}\hat{\sigma}^{12}\rangle_{\mathbf{q}=0}^{(2)(0)}$ for $\mathbf{q}=\mathbf{k}_{\parallel}$. The above expression depends parametrically on the pulse delay time $\Delta\tau$ via the Fourier transform of the probe pulse. This can be viewed as the third-order equivalent of Eq. (17). Note however, that at third order, the only effect of the saturation is the renormalization of the input field via the created population in the numerator of Eq. (21), while the lattice sum $\mathcal{S}_{\mathbf{q}}^{23}(\omega)$ describing the interactions among the particles remains unmodified. In fact, the third–order term can be viewed as the next term in the expansion of the previous linear‐response expression, with the fixed inversion replaced by its pump–induced, dynamically-modified value, while modification of the interaction strength would show up as a higher-order effect. The resulting extinction spectra, both linear in pump and probe, are shown in Figs. 6(b), (c). Importantly, we can see that the basic mechanism underlying SLR switching described in the previous section still remains valid even under the dynamic population assumption.
Figure 6: (a) Pump-probe spectroscopy of excitonic SLRs: Pump ($p$) and probe ($p^{\prime}$) pulses excite the emitters at time delay $\tau_{\Delta}$ (at normal incidence $k_{\ell,\parallel}=0$). The central frequencies of the pump and probe envelopes in frequency domain are tuned to the $\omega_{12}$- and $\omega_{23}$-transitions, respectively. For the pulses, we choose a width of $40\,\mathrm{fs}$ in time domain and identical pulse envelopes $f_{p}(t)=f_{p^{\prime}}(t)$. (b), (c) Linear extinction spectra in pump and probe field, at a time delay $\Delta\tau=50\,\mathrm{fs}$. The population created by the pump enables the $\omega_{23}$-transition, allowing the collective SLR resonance to emerge in the probe extinction spectrum. All other parameters are identical to Fig. 5.
VI Conclusions, outlook and extensions
@@
In the first application, we discussed how high-$Q$ SLRs can provide a pathway to bring molecular optomechanics in the resolved sideband regime and mediated the coupling between nanoparticle arrays and collective vibrational modes. In the second application, we demonstrated the versatility of SLRs in the nonlinear optical regime by considering arrays of multi-level emitters. Through optical pumping, the excitonic nonlinearity allowed for selective activation of specific SLR conditions corresponding to different electronic transitions. We have further shown that such switching can be observed and characterized in pump–probe experiments, for which we performed a perturbative expansion of the emitter dynamics up to third order, explicitly including population dynamics. To the best of our knowledge, this constitutes the first description of pump–probe spectroscopy in SLRs that does not rely on phenomenological models, but instead derives directly from a microscopic treatment of the emitter–field interactions.. This ability to dynamically control SLRs suggests potential uses in tunable photonic devices and optically driven switches, particularly in systems where reconfigurable optical properties are desirable.
For future endeavors, the quantized description of the electromagnetic field employed in our treatment could e.g., be used to investigate non-classical states of light such as squeezed light input, with possible applications in quantum-enhanced sensing. While so far, as a proof of principle, we have focused on extinction spectra, in principle we could also compute more complex material observables, like molecular populations through nonlinear optical spectroscopy, using mean-field approaches Li et al. (2018); Fowler-Wright et al. (2022); Reitz et al. (2025). Finally, while we have focused on a simplified treatment of the nanoparticles as harmonic oscillators in the dipole approximation, the approach can be extended to more complicated treatments, e.g., based on quasi-normal mode theory or higher-order multipoles Miwa and Schatz (2021).
VII Acknowledgments
@@
$\times\mathrm{e}^{-\mathrm{i}\mathbf{q}\cdot\mathbf{r}_{j}}\boldsymbol{\varepsilon}_{\boldsymbol{\mu}}\cdot\mathrm{Im}[\mathbf{G}(\mathbf{r}_{j},\mathbf{r}_{j^{\prime}},\omega_{\mathbf{k}})]\cdot\boldsymbol{\varepsilon}_{\boldsymbol{\mu}}\mathrm{e}^{\mathrm{i}\mathbf{q}^{\prime}\cdot\mathbf{r}_{j^{\prime}}}{\hat{A}}_{\mathbf{q}^{\prime}}(t^{\prime}).$
@@
$\times\sum_{\Lambda}\mathrm{e}^{-\mathrm{i}\mathbf{q}\cdot\mathbf{r}_{\Lambda}}\boldsymbol{\varepsilon}_{\boldsymbol{\mu}}\cdot\mathrm{Im}\left[\mathbf{G}(\mathbf{r}_{\Lambda},\omega_{\mathbf{k}})\right]\cdot\boldsymbol{\varepsilon}_{\boldsymbol{\mu}}{\hat{A}}_{\mathbf{q}}(\omega)\,\mathrm{e}^{\mathrm{i}\omega(t-t^{\prime})}.$ (S18)
@@
in terms of the single-particle radiative rate $\Gamma_{0}^{\mathrm{rad}}=(\mu_{0}^{2}\omega_{0}^{3})/(3\pi c^{3}\hbar\epsilon_{0})$. Here, the reduced lattice sum $\mathcal{S}_{\mathbf{q}}(\omega)$ now excludes the zero-displacement contribution and we made a Markovian assumption for the single-particle radiative rate only by evaluating it around $\omega_{0}\approx\omega$. The condition that needs to be fulfilled for SLRs is that both the real and the imaginary part of the expression in the brackets becomes (close to) zero. For numerical calculations, it is useful to reexpress the constants appearing in front of the lattice sum in terms of the single-particle radiative decay rate as
@@
with $\lambda(\omega)=2\pi c/\omega$. Equivalently, we remark that one could carry out the derivation entirely in real space and apply the lattice Fourier transform only at the end.
@@
Figure S1: Schematics of elimination procedure. Illustration of a nanoparticle array ($\hat{\boldsymbol{\mu}}_{j}$) and an array of dipolar emitters ($\hat{\mathbf{d}}_{j}$) which are commonly coupled to the background electromagnetic field modes $\hat{a}_{\mathbf{k},\lambda}$. Elimination of the electromagnetic field leads to effective interactions both within the arrays and between the nanoparticle and emitter array, which are encompassed in the corresponding lattice sums.
Let us now consider the additional coupling of the nanoparticle array to an (e.g., molecular) array of dipolar emitters. We use this as an instructive case to facilitate the derivation of the optomechanical interaction in the following section. We assume the individual emitters with resonance frequency $\omega_{d}$ located at positions $\mathbf{r}_{j}^{\mathrm{m}}=\mathbf{r}_{j}+\mathbf{r}_{\mathrm{m}}$ described by the dipole operator $\hat{\mathbf{d}}_{j}=\mathbf{d}_{0}\left(\hat{B}_{j}+\hat{B}_{j}^{\dagger}\right)$ with dipole moment $\mathbf{d}_{0}$ and annihilation/creation operators $\hat{B}_{j}^{\phantom{\dagger}},\hat{B}_{j^{\prime}}^{\dagger}]=\delta_{j,j^{\prime}}$ and typically $|\mathbf{d}_{0}|\ll|\boldsymbol{\mu}_{0}|$. We also assume that the emitter array has the same periodicity as the nanoparticle array. We start again by writing equations of motion for dipole operators as well as the electromagnetic field operators
@@
and replacing them in the equations for the dipole operators. The procedure is now largely identical to Sec. S1 and consists in going to Fourier space both in the time and spatial domains. Finally, one ends up with an algebraic set of equations for the dipole operators
@@
describing a sum over all interactions between nanoparticle and emitter array (see Fig. S1 for pictorial representation of lattice sums). The input noise terms $\hat{A}_{\mathbf{q}}^{\mathrm{in}}(\omega),\hat{B}_{\mathbf{q}}^{\mathrm{in}}(\omega)$ are defined analogously to Sec. S1. Assuming only the nanoparticles to be driven by a non-zero average input electric field, i.e., $\expectationvalue{\hat{B}_{\mathrm{in}}(\omega)}=0$, the dipole amplitude of the nanoparticle array can be expressed as
@@
$\hat{\mathcal{H}}_{\mathrm{OM}}=-\sum_{j}\hat{\mathbf{p}}_{j}^{R}\cdot\hat{\mathbf{E}}(\mathbf{r}_{j}^{\text{m}}).$ (S32)
@@
and the free energy of the vibrational modes is given by $\hat{\mathcal{H}}_{\mathrm{vib}}=\sum_{j}\omega_{\mathrm{vib}}\hat{b}_{j}^{\dagger}\hat{b}_{j}^{\phantom{\dagger}}$. With this, the optomechanical Hamiltonian can be written as
$\hat{\mathcal{H}}_{\mathrm{OM}}=-\sum_{j}\sum_{\mathbf{k},\lambda}\mathcal{E}_{\mathbf{k}}\left(\mathbf{p}_{j}\mathrm{e}^{-\mathrm{i}\omega_{\ell}t}+\mathbf{p}_{j}^{*}\mathrm{e}^{\mathrm{i}\omega_{\ell}t}\right)(\hat{b}_{j}^{\dagger}+\hat{b}_{j}^{\phantom{\dagger}})\left(\boldsymbol{\varepsilon}_{\mathbf{k},\lambda}\hat{a}_{\mathbf{k},\lambda}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{m}}+\boldsymbol{\varepsilon}_{\mathbf{k},\lambda}^{*}\hat{a}_{\mathbf{k},\lambda}^{\dagger}\mathrm{e}^{-\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{m}}\right).$ (S35)
@@
$\hat{\mathcal{H}}_{\mathrm{OM}}=\hbar\sum_{j}\sum_{\mathbf{k},\lambda}\left[g_{\mathbf{k},\lambda}^{\mathrm{OM}}\hat{a}_{\mathbf{k},\lambda}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{\mathrm{m}}}\mathrm{e}^{\mathrm{i}\omega_{\ell}t}+g_{\mathbf{k},\lambda}^{\mathrm{OM},*}\hat{a}_{\mathbf{k},\lambda}^{\dagger}\mathrm{e}^{-\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{\mathrm{m}}}\mathrm{e}^{-\mathrm{i}\omega_{\ell}t}\right](\hat{b}_{j}^{\dagger}+\hat{b}_{j}^{\phantom{\dagger}}),$ (S36)
@@
$\hat{\mathcal{H}}_{\mathrm{OM}}^{\mathrm{AS}}\approx\hbar\sum_{j}\sum_{\mathbf{k},\lambda}\left[g_{\mathbf{k},\lambda}^{\mathrm{OM}}\hat{a}_{\mathbf{k},\lambda}\hat{b}_{j}^{\dagger}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{\mathrm{m}}}\mathrm{e}^{\mathrm{i}\omega_{\ell}t}+g_{\mathbf{k},\lambda}^{\mathrm{OM},*}\hat{a}_{\mathbf{k},\lambda}^{\dagger}\hat{b}_{j}\mathrm{e}^{-\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{\mathrm{m}}}\mathrm{e}^{-\mathrm{i}\omega_{\ell}t}\right],$ (S37)
@@
$\hat{\mathcal{H}}_{\mathrm{OM}}^{\mathrm{S}}\approx\hbar\sum_{j}\sum_{\mathbf{k},\lambda}\left[g_{\mathbf{k},\lambda}^{\mathrm{OM}}\hat{a}_{\mathbf{k},\lambda}\hat{b}_{j}^{\phantom{\dagger}}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{\mathrm{m}}}\mathrm{e}^{\mathrm{i}\omega_{\ell}t}+g_{\mathbf{k},\lambda}^{\mathrm{OM},*}\hat{a}_{\mathbf{k},\lambda}^{\dagger}\hat{b}_{j}^{\dagger}\mathrm{e}^{-\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{\mathrm{m}}}\mathrm{e}^{-\mathrm{i}\omega_{\ell}t}\right],$ (S43)
@@
S4. Nonlinear polarizability of excitonic SLRs for static populations